# User-Creator Feature Dynamics in Recommender Systems with Dual Influence

### v1

@@ Abstract
Recommender systems present relevant contents to users and help content creators reach their target audience. The dual nature of these systems influences both users and creators: users’ preferences are affected by the items they are recommended, while creators are incentivized to alter their contents such that it is recommended more frequently. We define a model, called user-creator feature dynamics, to capture the dual influences of recommender systems. We prove that a recommender system with dual influence is guaranteed to polarize, causing diversity loss in the system. We then investigate, both theoretically and empirically, approaches for mitigating polarization and promoting diversity in recommender systems. Unexpectedly, we find that common diversity-promoting approaches do not work in the presence of dual influence, while relevancy-optimizing methods like top-$k$ recommendation can prevent polarization and improve diversity of the system.
Recommender Systems, Preference Dynamics, Dual Influence
@@ 1 Introduction
From restaurant selection, video watching, to apartment renting, recommender systems play a pivotal role across a plethora of real-world domains. These systems match users with items they like, and help creators (those producing the items) identify their target audiences. Nevertheless, behind such success, concerns have emerged regarding possible harmful outcomes of recommender systems, in particular, filter bubbles (Masrour et al., 2020; Aridor et al., 2020) and polarization (Santos et al., 2021) – outcomes with insufficient recommendation diversity and creation diversity. Recommendation diversity, meaning the diversity of the contents recommended to a user, is key to users’ engagement and retention on the platform. Meanwhile, creation diversity, meaning the variety of content created on the platform, is a determinant of the platform’s long-term health. In extreme cases, insufficient creation diversity can lead to consensus or polarization, where the latter can cause conflict and hatred, diminish people’s mutual understanding, and cause societal crises. Therefore, from both business and social responsibility perspectives, championing and improving diversity in recommender systems is equally important as optimizing recommendation relevance.
@@
There is increasing emphasis in academia and industry on investigating and improving the diversity of recommender systems, combating filter bubbles and polarization. Popular diversity-boosting approaches include applying post-processing procedures such as re-ranking (Carbonell & Goldstein, 1998; Ziegler et al., 2005) and setting diversity-aware objectives in addition to relevance maximization (Su et al., 2013; Zhang & Hurley, 2008; Hurley, 2013; Wilhelm et al., 2018; Cheng et al., 2017). These methods aim to increase the recommendation diversity for users. Assuming that the contents on the platform are static, these methods have been shown to bring diversity gain to the system.
@@
However, an important aspect that is overlooked in the aforementioned approaches is that: users and contents on a recommendation platform are not static entities – they can be influenced by the recommendation made by the system. In content creation platforms like YouTube, TikTok, and Twitter, recommendations naturally affect both content users and content creators. It is well known that the exposure to recommended items can shift a user’s preference (Jiang et al., 2019; Dean & Morgenstern, 2022). On the other hand, the creators have the incentive to change their creation styles constantly to attract their audience better (and to make more profits from the platform) (Eilat & Rosenfeld, 2023; Hron et al., 2023; Jagadeesan et al., 2024). While the effects of recommendation on either users or creators have been investigated separately, no previous work considers both effects to our knowledge. The dual influence of recommendation on users and creators causes complicated dynamics where users and creators interact and their preferences evolve together. Such evolution might exacerbate filter bubbles and polarization. Whether the aforementioned diversity-boosting approaches still work in a dynamic environment with dual influence is questionable.
@@
The first contribution of our work is to define a novel, natural dynamics model that captures the dual influence of a recommender system on users and creators, which we call user-creator feature dynamics (Section 2). We leverage the users’ and items’/creators’ embedding vectors to represent their preferences and creation styles, and use cosine similarity to characterize the relevance of creations and users’ interests (which is common in the recommender system literature and practice). This model allows us to formally reason about the impact of various design choices on the diversity of a recommender system with dual influence.
@@
Our second contribution is to demonstrate that, under realistic conditions, the user-creator feature dynamics of any recommender system with dual influence is guaranteed to converge to polarization (Section 3), i.e., the preferences of users and the contents of creators will be tightly clustered into two opposite groups (bringing the diversity of the system to nearly zero). We demonstrate that this phenomenon still occurs even after applying diversity-boosting approaches to the recommender system.
@@
Then, (in Section 4) we investigate some real-world designs of recommendation algorithms in order to look for techniques that mitigate polarization. Interestingly, we find that some common efficiency-improving methods, such as top-$k$ recommendation, can both prevent the system from polarization and improve the creation diversity. We also provide empirical results on both synthetic and real-world (MovieLens) data that support our theoretical claims; see Section 5 for a summary. Section 6 concludes.
Additional Related Works
The most related works have been mentioned above. App. A offers additional discussions on three strands of literature: diversity in recommender systems, opinion dynamics, and performative prediction.
@@ 2 Model: User-Creator Feature Dynamics
$p_{ij}^{t}=\mathrm{softmax}(\bm{u}_{j}^{t},\bm{V}^{t};\beta)=\tfrac{\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}.$ (2)
@@
Here, $\eta_{u}>0$ is a parameter controlling the rate of update, $f(\bm{v}_{i},\bm{u}_{j})$ is a function that quantifies the impact of creator $i$’s content on user $j$, and $\mathrm{\mathcal{P}}(\bm{x})=\frac{\bm{x}}{\|\bm{x}\|_{2}}$ is the projection onto the unit sphere. Our user update model generalizes (Dean & Morgenstern, 2022).
@@
where $\eta_{c}>0$ is an update rate, and $g(\bm{u}_{j},\bm{v}_{i})$ is a function that quantifies the impact of user $j$ on creator $i$. Our creator update model generalizes (Eilat & Rosenfeld, 2023).
@@
Our results apply to any impact functions $f$ and $g$ that satisfy the following natural assumptions. First, $f(\bm{v}_{i},\bm{u}_{j})$ and the inner product $\langle\bm{v}_{i},\bm{u}_{j}\rangle$ have the same sign: $f(\bm{v}_{i},\bm{u}_{j})\text{ is }{\tiny\begin{cases}>0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle>0\\ <0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle<0\\ =0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle=0.\end{cases}}$ This means that if a user likes the content ($\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle>0$), then the user vector $\bm{u}_{j}^{t}$ will be updated towards the direction of the creator vector $\bm{v}_{j}^{t}$. If the user dislikes the content ($\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle<0$), then the user vector $\bm{u}_{j}^{t}$ will move away from $\bm{v}_{j}^{t}$. Further assume upper and lower bounds on $|f|$:
$|f(\bm{v}_{i},\bm{u}_{j})|\leq 1,\quad~{}|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}>0\text{ when $\langle\bm{v}_{i},\bm{u}_{j}\rangle\neq 0$}.$
@@
For $g$, likewise assume that its sign is the same as $\langle\bm{u}_{j},\bm{v}_{i}\rangle$: $g(\bm{u}_{j},\bm{v}_{i})\text{ is }{\tiny\begin{cases}>0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle>0\\ <0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle<0\\ =0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle=0.\end{cases}}$ Intuitively, this captures the incentive of a creator who tries to maximize the average ratings from users who are recommended their items. On video platforms for example, if the creators are rewarded based on the average rating of their videos, they will try to reinforce their creation style based on the users who give positive feedback ($\langle\bm{u}_{j},\bm{v}_{i}\rangle>0$) so that their creations are more likely to be recommended to those users. Meanwhile, they will also change their creation style based on negative feedback ($\langle\bm{u}_{j},\bm{v}_{i}\rangle<0$), but in the opposite direction of the negative-feedback users’ interests, so that their creations are less likely to be recommended to those users. A particular example of $g$ is the sign function $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)\in\{-1,0,1\}$. We will focus on the sign function $g$ to simplify the theoretical presentation. All our other results hold for general $g$ functions.
@@ 3 Unavoidable Polarization
Having defined the user-creator feature dynamics in a recommender system with dual influence, we theoretically study how such dynamics evolve. Our main result is: if every creator can be recommended to every user with some non-zero probability, then the dynamics must eventually polarize.
@@
  • $R$-bi-polarization if there exists a vector $\bm{c}\in{\mathbb{R}}^{d}$ such that every feature vector is $R$-close to $+\bm{c}$ or $-\bm{c}$: $\forall\bm{u}_{j}^{t}$, $\|\bm{u}_{j}^{t}-\bm{c}\|_{2}\leq R$ or $\|\bm{u}_{j}^{t}+\bm{c}\|_{2}\leq R$ holds, and $\forall\bm{v}_{i}^{t}$, $\|\bm{v}_{i}^{t}-\bm{c}\|_{2}\leq R$ or $\|\bm{v}_{i}^{t}+\bm{c}\|_{2}\leq R$ holds.
Consensus is any state where all users and creators have similar feature vectors (up to difference $R$), implying that they have similar preferences. Bi-polarization is any state where all users and creators are clustered into two groups with exactly opposite features (e.g., Republicans vs Democrats).
@@
Bi-polarization states are absorbing: namely, once the dynamics reaches $R$-bi-polarization with some $R\in[0,1]$, it will satisfy $R$-bi-polarization forever. The same holds for consensus.
A natural property of a recommender system is that every creator can be recommended to every user with some non-zero probability: $p_{ij}^{t}\geq p_{0}>0$. This is satisfied by the softmax function, which is a rough model of real-world recommendations (Covington et al., 2016): $p_{ij}^{t}=\tfrac{\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}\geq\tfrac{\exp(-\beta)}{n\exp(\beta)}=p_{0}>0$. Moreover, many large-scale real-world recomendation systems (e.g., Yahoo! (Li et al., 2010) and Kuaishou (Gao et al., 2022)) intentionally insert small random traffic to improve recommendation diversity or to explore users’ interests (Judith Möller & van Es, 2018; Yang et al., 2018), which will cause all recommendation probabilities to be non-zero. We show that, however, a recommender system satisfying $p_{ij}^{t}\geq p_{0}>0$ must converge to polarization, under some additional conditions on the users’ and creators’ update rates:
@@
Suppose $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$, the update rates $\eta_{u}<\frac{1}{2}$ and $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$, and the recommendation probability $p_{ij}^{t}\geq p_{0}>0,\forall i,j,t$. Then, from almost all initial states, the dynamics $(\bm{U}^{t},\bm{V}^{t})$ will eventually reach $R$-consensus or $R$-bi-polarization for any $R>0$.
In other words, if the users’ and creators’ updates are not too fast and the recommendation probabilities are non-zero, then all users and creators will eventually converge to at most two clusters (regardless of the feature dimension $d$). Since creators in one cluster produce similar contents, users in such a polarized system can never receive diverse recommendations. This means that the naïve attempt of imposing $p_{ij}^{t}\geq p_{0}>0$ cannot improve the diversity of a recommender system with dual influence. The conditions on the update rates $\eta_{u},\eta_{c}$ are only assumed to simplify the proof of Theorem 3.3. Our experiments (in Appendix B) show that polarization still occurs even without those conditions.

### v2

# User-Creator Feature Polarization in Recommender Systems with Dual Influence

@@ Abstract
Recommender systems serve the dual purpose of presenting relevant content to users and helping content creators reach their target audience. The dual nature of these systems naturally influences both users and creators: users’ preferences are affected by the items they are recommended, while creators may be incentivized to alter their content to attract more users. We define a model, called user-creator feature dynamics, to capture the dual influence of recommender systems. We prove that a recommender system with dual influence is guaranteed to polarize, causing diversity loss in the system. We then investigate, both theoretically and empirically, approaches for mitigating polarization and promoting diversity in recommender systems. Unexpectedly, we find that common diversity-promoting approaches do not work in the presence of dual influence, while relevancy-optimizing methods like top-$k$ truncation can prevent polarization and improve diversity of the system.
@@ 1 Introduction
From restaurant selection, video watching, to apartment renting, recommender systems play a pivotal role across a plethora of real-world domains. These systems match users with content they like, and help creators (those producing the content) identify their target audiences. Nevertheless, behind such success, concerns have emerged regarding possible harmful outcomes of recommender systems, in particular, filter bubbles [32, 5] and polarization [36] – outcomes with insufficient recommendation diversity and creation diversity. Recommendation diversity, meaning the diversity of the contents recommended to a user, is key to users’ engagement and retention on the platform. Meanwhile, creation diversity, meaning the variety of content created on the platform, is a determinant of the platform’s long-term health. In extreme cases, insufficient creation diversity can lead to consensus or polarization, where the latter can cause conflict and hatred, diminish people’s mutual understanding, and cause societal crises. Therefore, from both business and social responsibility perspectives, championing and improving diversity in recommender systems is equally important as optimizing recommendation relevancy.
@@
There is increasing emphasis in academia and industry on investigating and improving the diversity of recommender systems, combating filter bubbles and polarization. Popular diversity-boosting approaches include applying post-processing procedures such as re-ranking [11, 47] and setting diversity-aware objectives in addition to relevance maximization [38, 44, 22, 39, 12]. These methods aim to increase the recommendation diversity for users. Assuming that the contents on the platform are static, these methods have been shown to bring diversity gain to the system.
@@
However, an important aspect that is overlooked in the aforementioned approaches is that: users and contents on a recommendation platform are not static entities – they can be influenced by the recommendation made by the system. In content creation platforms like YouTube, TikTok, and Twitter, recommendations naturally affect both content users and content creators. It is well known that the exposure to recommended items can shift a user’s preference [24, 26, 14]. On the other hand, the creators have the incentive to change their creation styles constantly to attract their audience better (and to make more profits from the platform) [15, 20, 23]. While the effects of recommendation on either users or creators have been investigated separately, to our knowledge no previous work considers both effects. The dual influence of recommendation on users and creators causes complicated dynamics where users and creators interact and their preferences evolve together. Such evolution might exacerbate filter bubble and polarization effects. Whether the aforementioned diversity-boosting approaches still work in a dynamic environment with dual influence is questionable.
Our contributions
@@
The first contribution of our work is to define a novel, natural dynamics model that captures the dual influence of a recommender system on users and creators, which we call user-creator feature dynamics (Section 2). We leverage the users’ and items’/creators’ embedding vectors to represent their preferences and creation styles, and use cosine similarity to characterize the relevance of creations and users’ interests (which is common in the recommender system literature and practice). This model allows us to formally reason about the impact of various design choices on the long-term diversity of a recommender system with dual influence.
@@
Our second contribution is to demonstrate that, under realistic conditions, the user-creator feature dynamics of any recommender system with dual influence must unavoidably converge to polarization (Section 3), i.e., the preferences of users and the contents of creators will become tightly clustered into two opposite groups, significantly reducing the diversity of the system. We demonstrate that this phenomenon still occurs even after applying diversity-boosting interventions to the system.
@@
Then, (in Section 4) we investigate some real-world designs of recommendation algorithms in order to look for techniques that mitigate polarization. Interestingly, we find that some common efficiency-improving methods, such as top-$k$ truncation, can both prevent the system from polarization and improve the creation diversity. We also provide empirical results (Section 5) on both synthetic and real-world (MovieLens) data. As predicted by our theory, we find that systems with dual influence more easily converge to polarization under diversity-boosting designs, while efficiency-oriented and relevance-optimizing designs can in fact improve the long-term diversity of the system. This could explain why polarization does not always happen in reality. Section 6 concludes.
1.1 Related Work
Diversity in recommendations
Diversity, filter bubbles, and polarization in recommendations have been important research topics in recent years, and they are closely related concepts with different focuses. On the one hand, filter bubbles are frequently defined as decreasing recommendation diversity over time [5], which describes both the process and the outcome of insufficiently diverse recommendations. On the other hand, polarization describes the negative outcome of insufficient mutual understanding between people [36]. In content platforms, an example of polarization is people creating content with strong agreement or disagreement with other content under the same topic, e.g., political opinions. To combat these negative outcomes, previous works propose diversity-boosting approaches including re-ranking [11, 47] and diversity-aware objective optimization [38, 44, 22, 39, 12, 45]. Despite having positive effects in situations where user preferences and creation styles are fixed, these approaches overlooked the dynamic nature of recommender systems and our work shows that certain approaches will make long-term outcomes worse under the dual influence.
Opinion dynamics
Opinion dynamics study the effect of people exchanging opinions with others on social networks [37, 17, 29, 4]. Our model of a recommender system with dual influence on users and creators resembles a bipartite social network, and our conclusion that the system converges to polarization is conceptually similar to people reaching consensus on social networks [1, 10, 31, 46]. However, the technique we use to prove our conclusion (absorbing Markov chain) significantly differs from the main technique (stability of ODE) in the mentioned works.
Performative effects of recommender systems
The phenomenon that predictive systems like recommender systems can impact the individuals interacting with those systems (e.g., users and creators) is related to the literature of performative prediction [34, 18]. These impacts can be direct, such as individuals ostensibly modifying their features in order to obtain more desirable outcomes [27]. Prior works on the performative effects of recommender systems (e.g., [7, 24, 14, 41, 15, 42, 35, 20, 3, 43, 2, 23]) only consider one-sided impact, either on users or on creators. Differing from them, our work studies two-sided impacts, i.e., on both users and creators. We provide a table to compare our work with previous works in Appendix A.
@@ 2 Model: User-Creator Feature Dynamics
$p_{ij}^{t}=\mathrm{softmax}(\bm{u}_{j}^{t},\bm{V}^{t};\beta)=\frac{\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}.$ (2)
@@
Here, $\eta_{u}\in[0,1]$ is a parameter controlling the rate of update, $f(\bm{v}_{i},\bm{u}_{j})$ is a function that quantifies the impact of creator $i$’s content on user $j$ (discussed in detail later), and $\mathrm{\mathcal{P}}(\bm{x})=\frac{\bm{x}}{\|\bm{x}\|_{2}}$ is the projection back onto the unit sphere. Our user update model generalizes [14], which considers $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}\langle\bm{v}_{i_{j}^{t}}^{t},\bm{u}_{j}^{t}\rangle\bm{v}_{i_{j}^{t}}^{t})$, by replacing the inner product with a general function $f$.
@@
where $\eta_{c}\in[0,1]$ is a parameter controlling the rate of update, and $g(\bm{u}_{j},\bm{v}_{i})$ is a function that quantifies the impact of user $j$ on creator $i$.
@@
Our results apply to any impact functions $f$ and $g$ that satisfy the following natural assumptions. First, $f(\bm{v}_{i},\bm{u}_{j})$ and the inner product $\langle\bm{v}_{i},\bm{u}_{j}\rangle$ have the same sign: $f(\bm{v}_{i},\bm{u}_{j})\text{ is }{\tiny\begin{cases}>0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle>0\\ <0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle<0\\ =0&\text{ if }\langle\bm{v}_{i},\bm{u}_{j}\rangle=0.\end{cases}}$ This means that if a user likes the content ($\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle>0$), then the user vector $\bm{u}_{j}^{t}$ will be updated towards the direction of the creator vector $\bm{v}_{j}^{t}$. If the user dislikes the content ($\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle<0$), then the user vector $\bm{u}_{j}^{t}$ will move away from $\bm{v}_{j}^{t}$. Such “biased assimilation” user behavior is well documented in the literature [14]. Further, we assume upper and lower bounds on $|f|$:
$|f(\bm{v}_{i},\bm{u}_{j})|\leq 1,\quad\quad|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}>0\text{ whenever $\langle\bm{v}_{i},\bm{u}_{j}\rangle\neq 0$}.$
@@
For $g$, likewise assume that its sign is the same as $\langle\bm{u}_{j},\bm{v}_{i}\rangle$: $g(\bm{u}_{j},\bm{v}_{i})\text{ is }{\tiny\begin{cases}>0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle>0\\ <0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle<0\\ =0&\text{ if }\langle\bm{u}_{j},\bm{v}_{i}\rangle=0.\end{cases}}$ Intuitively, this captures the incentive of a creator who aims to maximize the average ratings from users who are recommended their items. On video platforms for example, if the creators are rewarded based on the average rating of their videos, they will try to reinforce their creation styles based on the users who give positive feedback ($\langle\bm{u}_{j},\bm{v}_{i}\rangle>0$) so that their creations are more likely to be recommended to those users. Meanwhile, the creators will also change their creation styles based on negative feedback ($\langle\bm{u}_{j},\bm{v}_{i}\rangle<0$), but in the opposite direction of the negative-feedback users’ interests, so that their creations are less likely to be recommended to those users. Taking both scenarios into account, the creator moves towards the weighted average of user preferences $\sum_{j\in J_{i}^{t}}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t}$, which is captured by our update rule (4). A particular example of $g$ is the sign function $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)\in\{-1,0,1\}$. We will only consider the sign function $g$ in order to simplify the theoretical presentation. We believe that all our results can be generalized to other $g$ functions satisfying similar conditions as $f$; the details are left as future work.
@@ 3 Unavoidable Polarization
Having defined the user-creator feature dynamics in a recommender system with dual influence, we now theoretically study how such dynamics evolve. Our main result is: if every creator can be recommended to every user with some non-zero probability, then the dynamics must eventually polarize.
@@
  • $R$-bi-polarization if there exists a vector $\bm{c}\in{\mathbb{R}}^{d}$ such that every feature vector is $R$-close to $+\bm{c}$ or $-\bm{c}$: $\forall\bm{u}_{j}^{t}$, $\|\bm{u}_{j}^{t}-\bm{c}\|_{2}\leq R$ or $\|\bm{u}_{j}^{t}+\bm{c}\|_{2}\leq R$, and $\forall\bm{v}_{i}^{t}$, $\|\bm{v}_{i}^{t}-\bm{c}\|_{2}\leq R$ or $\|\bm{v}_{i}^{t}+\bm{c}\|_{2}\leq R$.
The dynamics is said to reach $(R,\bm{c})$-consensus (or $(R,\bm{c})$-bi-polarization) if the dynamics reaches $R$-consensus (or $R$-bi-polarization) with the vector $\bm{c}$.
Consensus is any state where all users and creators have similar feature vectors (with maximum difference $R$), implying that they have similar interests or preferences. Bi-polarization is any state where all users and creators are clustered into two groups with exactly opposite features (e.g., Republicans vs Democrats). Mathematically, consensus is a special case of bi-polarization.
@@
Bi-polarization states are absorbing: once the dynamics reaches $(R,\bm{c})$-bi-polarization with some $R\in[0,1]$ and $\bm{c}\in\mathbb{S}^{d-1}$, it will satisfy $(R,\bm{c})$-bi-polarization forever. The same holds for consensus.
A natural property of a recommender system is that every creator can be recommended to every user with some non-zero probability: $p_{ij}^{t}\geq p_{0}>0$ with some constant $p_{0}$. This is satisfied by the softmax function, which is a rough model of real-world recommendation algorithms [13, 26]: $p_{ij}^{t}=\tfrac{\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle)}\geq\tfrac{\exp(-\beta)}{n\exp(\beta)}=p_{0}>0$. Moreover, many large-scale real-world recomendation systems (e.g., Yahoo! [28] and Kuaishou [16]) intentionally insert small random traffic attempting to improve recommendation diversity or explore users’ interests [25, 40], which will cause all recommendation probabilities to be non-zero. We show in Theorem 3.3 that, however, a recommender system satisfying $p_{ij}^{t}\geq p_{0}>0$ must converge to polarization, under some additional conditions on the users’ and creators’ update rates:
@@
Suppose $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$, the update rates $\eta_{c}\leq\frac{\eta_{u}L_{f}}{2}$ and $\eta_{u}<\frac{1}{2}$, and the recommendation probability $p_{ij}^{t}\geq p_{0}>0,\forall i,j,t$. Then, from almost all initial states, the dynamics $(\bm{U}^{t},\bm{V}^{t})$ will eventually reach $R$-consensus or $R$-bi-polarization for any $R>0$.
In other words, if the users’ and creators’ updates are not too fast and all recommendation probabilities are non-zero, then all users and creators will eventually converge to at most two clusters (regardless of the feature dimension $d$). Since creators in one cluster produce similar contents, users in such a polarized system can never receive diverse recommendations. This means that the naïve attempt of imposing $p_{ij}^{t}\geq p_{0}>0$ cannot improve the diversity of a recommender system with dual influence. The conditions on the update rates $\eta_{u},\eta_{c}$ are only assumed to simplify the proof of Theorem 3.3. Our experiments (in Section 5) will show that polarization still occurs even without those conditions.
Theorem 3.3 does not characterize the rate of convergence of the user-creator feature dynamics to polarization, which we leave as an open question.
The proof of Theorem 3.3 is an absorbing Markov chain argument. It uses the following lemma:
Lemma 3.4.
Proof of Theorem 3.3.
For any state $(\bm{U}^{t},\bm{V}^{t})$ in the state space, by Lemma 3.4 there exists a path $(\bm{U}^{t},\bm{V}^{t})\to\cdots\to(\bm{U}^{t+T},\bm{V}^{t+T})$ of length $T$ that leads to $R$-bi-polarization. Because every creator can be recommended to a user with probability at least $p_{0}$, each transition $(\bm{U}^{t^{\prime}},\bm{V}^{t^{\prime}})\to(\bm{U}^{t^{\prime}+1},\bm{V}^{t^{\prime}+1})$ happens with probability at least $p_{0}^{m}$. So, the path of length $T$ has probability at least $p_{0}^{mT}>0$, and the probability that the dynamics does not reach $R$-bi-polarization after $KT$ steps is at most $(1-p_{0}^{mT})^{K}$, which $\to 0$ as $K\to\infty$. Therefore, with probability $1$ the dynamics will reach $R$-bi-polarization eventually. ∎